OpenStudy (savannahbean36):

The area in square meters of a rectangular parking lot is x^2-95x+2100 the width in meters is x-60 what is the length of the parking lot in meters show your work? I'm having a hard time with this can someone help

OpenStudy (karim728):

factor out the eqUATION

OpenStudy (holsteremission):

\[\text{area of rectangle}=\text{length}\times\text{height}\]You're given two bits of info here, the total area and the width/length. Replacing what you know in the equation above and abbreviating what you don't with a variable, you can write the equation above as \[x^2-95x+2100=(x-60)h\]As long as \(x\neq60\) (which it isn't, because that would mean the length is \(0\)), you can divide both sides by \(x-60\). Then \[h=\frac{x^2-95x+2100}{x-60}\]where the height \(h\) is in meters. Can you continue from here?